OpenStudy (anonymous):

An isosceles trapezoid has base angles equal to 45 and bases of lengths 6 and 12. Find the area of the trapezoid.

OpenStudy (anonymous):

The height of the trapezoid is 3 since the shape is isosceles trapezoid and is 45-45-90 degree triangle, and to find its height, subtract the bases and then divide the answer by 2. (12-6)/2 6/2=3 Now find the area of a trapezoid Area = average of bases times height where b1, b2 are the lengths of each base and h is the height. Area = (b 1 + 2)/2 x h Area = (12 + 6)/2 x 3 Area = 9 x 3 Area = 27

OpenStudy (anonymous):

@thockenberry14 u there?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

i understand it now

OpenStudy (anonymous):

Area = (b 1 + 2)/2 x h i messed up in the area it should be Area = (b 1 + b2)/2 x h No problem ::)

OpenStudy (anonymous):

ok